### v1

@@ Abstract
To establish a self-sustained human presence in space and to explore deeper into the solar system, extensive research has been conducted on In-Situ Resource Utilization (ISRU) systems. Past studies have proposed and researched many technologies to produce oxygen from regolith, such as carbothermal reduction and water extraction from icy regolith, to utilize it for astronauts’ life support and as the propellant of space systems. However, determining the most promising technology remains challenging due to uncertainties in the lunar environment and processing methods. To better understand the lunar environment and ISRU operations, it is crucial to gather more information. Motivated by this need for information gathering, this paper proposes a new ISRU plant architecture integrating carbothermal reduction of dry regolith and water extraction from icy regolith. Two different hybrid plant architectures integrating both technologies (1) in parallel and (2) in series are examined. The former involves mining and processing in both a Permanently Shadowed Region (PSR) and a peak of eternal light in parallel, while the latter solely mines in a PSR. In this series hybrid architecture, the dry regolith tailings from water extraction are further processed by carbothermal reduction. This paper conducts a comparative analysis of the landed mass and required power of each plant architecture utilizing subsystem-level models. Furthermore, based on uncertain parameters such as resource content in regolith, the potential performance range of each plant was discovered through Monte Carlo simulations. The result indicates the benefit of the series hybrid architecture in terms of regolith excavation rate, while its mass cost seems the highest among the studied architectures.
@@
{highlights}
Research highlight 1
Research highlight 2
@@ 1 Introduction
In-Situ Resource Utilization (ISRU), also known as Space Resource Utilization (SRU), aims to process and utilize local extraterrestrial resources to enable long-term human presence beyond low Earth orbit. Resources on our nearest planetary bodies, i.e., the Moon and Mars, include water in the form of ice or hydrated minerals [1, 2, 3], atmospheric carbon dioxide [2], and regolith, which is a source of oxygen and metals [1, 3]. These resources can be processed by a variety of thermo- and electrochemical techniques [2, 4] to produce life support consumables (e.g., H2O, O2), rocket fuel propellant (e.g., CH4, H2, O2), or infrastructure and spare parts (e.g., regolith, metals). Among various ISRU activities, lunar ISRU has been increasingly gathering a large amount of attention in response to the concurrent Artemis program and other initiatives of going back to the Moon.
@@
Two different types of hybrid lunar ISRU architectures are proposed. In the first architecture, carbothermal reduction and direct water extraction from icy regolith occur simultaneously in parallel. In the other architecture, these two processes are integrated in series, i.e., the remaining dry regolith after water extraction is further processed through carbothermal reduction. The Concept of Operations (ConOps) suitable for each architecture is carefully defined. Although these proposed hybrid architectures add complexity to the overall system, they can mitigate the risk of failure by avoiding reliance on a single extraction technology.
This paper further compares the mass and power budgets, along with other system performance indicators, of the proposed hybrid plant architecture with single technology architectures: carbothermal reduction of dry regolith and direct water extraction from icy regolith. A holistic and multidisciplinary approach that models end-to-end ISRU production plants is utilized to understand the benefits and drawbacks of specific technologies and processing techniques [28].
@@
The rest of the paper is organized as follows: Section 2 summarizes past studies in hybrid ISRU architectures, mass and power estimation of ISRU systems, and uncertainty related to lunar ISRU highlighting the research gap addressed by this paper. Section 3 introduces the hybrid ISRU production plant concept, including assumptions made and ConOps. In Section 4, the modeling approach of this work for the mass and power estimations is explained. The results of the mass and power estimations are discussed further in Section 5. Section 6 introduces the intrinsic uncertainty of the lunar environment and ISRU systems, and its effects on the performance of each ISRU plant operation. Section 7 discusses the potential benefits and risks of each plant design as well as critical parameters affecting the performance significantly. Finally, concluding remarks are outlined in Section 8.
@@ 2.1 Hybrid ISRU architecture
Chen et al. [27] also showed the benefit of employing both hydrogen reduction and water extraction from Martian soil through the optimization of technology selection in terms of crewed Mars mission cost. Even though this study showed the economic benefit, they also noted the potential increase in the complexity of its development, deployment, and operation.
@@ 2.2 Mass and power estimation for ISRU
As an example of relatively low-fidelity modeling, Chen et al. [27] developed a database of subsystem-level specific masses and powers based on literature, such as Ref. [33], and integrated it into an Earth-Moon-Mars space resources logistics network. This approach analyzed the effect of ISRU deployment location (i.e., the Moon or a near-Earth object) regarding system mass and monetary mission cost.
While these models can capture the trend in mass and power for different architectures, they often overlook the detailed ConOps, potentially missing some key subsystems and components to be considered. Recently, more detailed models considering ConOps carefully have been developed. Linne et al. [7] modeled in detail the extraction of oxygen via carbothermal reduction of dry regolith to identify the feasibility of deploying such a plant considering the capacity of a lunar lander. Guerrero-Gonzalez and Zabel [32] also accounted for metal processing by comparing three different ISRU techniques, hydrogen reduction, molten regolith electrolysis, and molten salt electrolysis, and estimating the mass and required power of each subsystem. Kiewiet et al. [22] modeled water extraction from icy regolith via three different heating methods and conducted a trade-off study. Kleinhenz and Paz [23] compared the system mass and power of two different architectures: the extraction of both H2 and O2 from water ice in the lunar southern polar region and solely O2 based on the work developed by Linne et al. [7].
@@ 2.3 Recognizing uncertainty in ISRU
None of these studies have compared different ISRU technologies considering the effect of the uncertain lunar environment and ISRU operations. Due to the large difference between dry and icy regolith processing, potential benefits and risks can be easily overlooked when ignoring the uncertainty. Inspired by Ref. [30], this paper models the distributions of uncertain parameters and examines their effects on plant performance indicators such as regolith excavation rate.
@@ 3 Hybrid lunar ISRU plant concept
This section first proposes two different hybrid lunar ISRU architectures. Technologies selected for further analyses and the detailed ConOps are explained in Sec. 3.2 and 3.3, respectively.
@@ 3.1 Overview and assumptions
Two different types of hybrid production plants are considered in this paper. The first architecture considered is a Parallel Hybrid plant that excavates and extracts resources from both PEL and PSR sites (Fig. 1(a)). In the PEL, dry regolith is excavated and processed to extract oxygen. Simultaneously in the PSR, water is extracted from icy regolith. The extracted water is transported from the PSR to the PEL by a mobile water transportation solution, where the water is then electrolyzed into O2 and H2. This concept is based on the water extraction plant design proposed by Kleinhenz and Paz [23]. Thermochemical extraction processes, such as carbothermal or hydrogen reduction of dry regolith, produce water as an intermediate product, and therefore, water electrolysis and storage of oxygen and hydrogen can be shared between dry and icy regolith processing.
The second architecture considered is a Series Hybrid, where dry regolith tailings from the water extraction of icy regolith are further processed directly in the PSR (Fig. 1(b)). This architecture can potentially maximize the oxygen and hydrogen yield from the regolith feedstock by utilizing both water ice and lunar regolith from excavated feedstock. Similar to the Parallel Hybrid architecture, the extracted and produced water is transported to a PEL for further processing. The detailed ConOps of both hybrid plants are further described in Section 3.3.
(a) Parallel Hybrid architecture. Dry regolith and icy regolith are processed simultaneously in different sites.
(b) Series Hybrid architecture. Dry regolith from water extraction process is further processed in a PSR.
Figure 1: Hybrid ISRU production plant architectures.
@@
Although identifying a specific site within the lunar South Pole was out of the scope of this work, previous studies [23, 42, 43] have identified traverse distances of around 0.8-8.5 km between potential human landing system sites, permanently shadowed craters where acceptable water ice deposits might be found, and ridges with prolonged sunlight availability. A preliminary analysis demonstrated that mass and power budgets were not sensitive to changes in the traverse distance. Therefore, an average traverse distance of 5 km between PSRs and PELs is employed in this study as depicted in Fig. 1.
As a baseline case scenario, the ISRU plant is assumed to produce 10 tons of O2 and 1.25 tons of H2 per year. This 8:1 ratio corresponds to the stoichiometric mass ratio of water to maximize the use of local resources for various purposes. It is worth noting that, for hydrolox lunar ascent vehicle propellant, the fuel mixture mass ratio is generally close to 6:1. Some past studies employed this 6:1 ratio or a similar ratio as their required products ratio allowing an excess of oxygen [23, 44].
The integrated hybrid production plant model in this paper includes all the necessary ISRU subsystems, from regolith excavation to cryo-storage of the produced O2 and H2 (see Section 4 for a detailed explanation of the models), except for power, thermal control (i.e., radiators), and water purification subsystems. The infrastructure and operations to further utilize these products are not addressed in this work.
@@
Table 1 collects the global assumptions considered in this study, including the radiative outer space and average surface temperatures for each site, which are used for the heat loss and radiative heat exchange calculations, and the dry and icy regolith density dependencies on depth and water ice content. Individual assumptions are described for each model in Section 4. Moreover, Section 6 discusses the values of the following uncertain parameters considered in this study: feedstock particle size, silicate and water ice content in the regolith, water extraction efficiency, and ISRU plant operational availability.
@@
[…]
1 z is the regolith depth in meters, with z = 0 m at the lunar surface. Based on Hayne et al. [46].
2 The icy regolith density is calculated assuming that water ice mass fraction (wt%ice) accumulates uniformly within the regolith pores.
@@ 3.2 Technology selection
Over the last decades, numerous strategies for extracting oxygen from dry lunar regolith have been proposed in the literature [25, 4, 47]. From these strategies, ilmenite reduction by hydrogen followed by the electrolysis of water and carbon-based reductants (e.g., CO, CH4, C) has been considered as feasible and viable options [47]. Carbothermal reduction of partially molten regolith, with water electrolysis and CH4-reforming, the direct electrochemical reduction of molten regolith, the electrolysis in a bath of molten salts, or vacuum thermal decomposition of lunar soils, have also been extensively researched [4].
Carbothermal Reduction (CR) of partially molten regolith is chosen as O2 extraction technology in this study due to its high yield, as it can reduce most regolith minerals, including the most prevalent silicates. Hydrogen reduction is excluded from further consideration since hydrogen is mostly limited to reducing iron-bearing minerals such as ilmenite, which are expected to only present in about 0.5 wt.of the lunar Sothern polar highland regolith [48]. As demonstrated by Guerrero-Gonzalez and Zabel [32], the low ilmenite content significantly increases the total hardware mass of the plant, highlighting its ineffectiveness as an oxygen extraction technology in these polar regions.
@@ 3.3.1 Parallel Hybrid architecture
Six systems are involved in the overall concept of operations of the Parallel Hybrid ISRU production architecture: (1) a dry-regolith mining (excavation and disposal) system, (2) a dry-regolith processing plant, (3) an H2/O2 production and storage plant, (4) an icy-regolith mining (excavation and disposal) system, (5) an icy-regolith processing plant, and (6) a water transportation system between both ISRU plants, as depicted in Fig. 1(a).
At the PEL, an excavator delivers dry regolith to the processing plant to extract oxygen. To avoid disturbances, the excavation site is located 100 m away from the ISRU facility [7]. After delivering a fresh regolith batch, the same vehicle is loaded with ISRU-produced tailings and drives first to a disposal site, also located 100 m away, before going back to the excavation site. The dry-regolith excavator recharges at the dry-regolith processing plant.
At the PSR, a similar scenario is defined. Another excavator delivers icy regolith to the icy-regolith processing plant, collects dry processed regolith, and dumps it at a disposal site before going back to the icy-regolith excavation site. The three locations are also positioned 100 m away from each other. The number of excavators for the dry- and icy-regolith is calculated based on the processing times, travel times, and regolith throughputs needed. The icy-regolith excavator only recharges at the icy-regolith processing plant without leaving the PSR.

### v2

@@ Abstract
To establish a self-sustained human presence in space and to explore deeper into the solar system, extensive research has been conducted on In-Situ Resource Utilization (ISRU) systems. Past studies have proposed and researched many technologies to produce oxygen from regolith, such as carbothermal reduction and water extraction from icy regolith, to utilize it for astronauts’ life support and as the propellant of space systems. However, determining the most promising technology remains challenging due to uncertainties in the lunar environment and processing methods. To better understand the lunar environment and ISRU operations, it is crucial to gather more information. Motivated by this need for information gathering, this paper proposes a new ISRU plant architecture integrating carbothermal reduction of dry regolith and water extraction from icy regolith. Two different hybrid plant architectures integrating both technologies (1) in parallel and (2) in series are examined. The former involves mining and processing in both a Permanently Shadowed Region (PSR) and a peak of eternal light in parallel, while the latter solely mines in a PSR. In this series hybrid architecture, the dry regolith tailings from water extraction are further processed by carbothermal reduction. This paper conducts a comparative analysis of the landed mass and required power of each plant architecture utilizing subsystem-level models. Furthermore, based on uncertain parameters such as resource content in regolith, the potential performance range of each plant was discovered through Monte Carlo simulations. The result indicates the benefit of the series hybrid architecture in terms of regolith excavation rate and power consumption, while its mass cost seems the highest among the studied architectures.
@@
Abbreviations: CHX, condensing heat exchanger; ConOps, concept of operations; COTS, commercial- off-the-shelf; CR, carbothermal reduction; ISRU; in-situ resource utilization; PEL, peak of eternal light; PEM, proton exchange membrane; PH, parallel hybrid; PSR, permanently shadowed region; LH2, liquid hydrogen; LOX, liquid oxygen; SH, series hybrid; SOXE, solid oxide electrolysis; WE, water extraction
@@ 1 Introduction
In-Situ Resource Utilization (ISRU), also known as space resource utilization, aims to process and utilize local extraterrestrial resources to enable long-term human presence beyond low Earth orbit. Resources on our nearest planetary bodies, i.e., the Moon and Mars, include water in the form of ice or hydrated minerals [1, 2, 3], atmospheric carbon dioxide [2], and regolith, which is a source of oxygen and metals [1, 3]. These resources can be processed by a variety of thermo- and electrochemical techniques [2, 4] to produce life support consumables (e.g., H2O, O2), rocket fuel propellant (e.g., CH4, H2, O2), or infrastructure and spare parts (e.g., regolith, metals). Among various ISRU activities, lunar ISRU has been increasingly gathering a large amount of attention in response to the concurrent Artemis program and other initiatives of going back to the Moon.
@@
Two different types of hybrid lunar ISRU architectures are proposed. In the first architecture, Carbothermal Reduction (CR) and direct Water Extraction (WE) from icy regolith occur simultaneously in parallel. In the other architecture, these two processes are integrated in series, i.e., the remaining dry regolith after WE is further processed through CR. The Concept of Operations (ConOps) suitable for each architecture is carefully defined. Although these proposed hybrid architectures add complexity to the overall system, they can mitigate the risk of failure by avoiding reliance on a single extraction technology.
This paper further compares the mass and power budgets, along with other system performance indicators, of the proposed hybrid plant architecture with single technology architectures: CR of dry regolith and direct WE from icy regolith. A holistic and multidisciplinary approach that models end-to-end ISRU production plants is utilized to understand the benefits and drawbacks of specific technologies and processing techniques [28].
@@
The rest of the paper is organized as follows: Section 2 summarizes past studies in hybrid ISRU architectures, mass and power estimation of ISRU systems, and uncertainty related to lunar ISRU highlighting the research gap addressed by this paper. Section 3 introduces the hybrid ISRU production plant concept, including assumptions made and ConOps. In Section 4, the modeling approach of this work for the mass and power estimations is explained. The results of the mass and power estimations are discussed further in Section 5. Section 6 introduces the intrinsic uncertainty of the lunar environment and ISRU systems, and its effects on the performance of each ISRU plant operation. Section 7 discusses the potential benefits and risks of each plant design as well as limitations of this study and directions for future work. Finally, concluding remarks are outlined in Section 8.
@@ 2.1 Hybrid ISRU architecture
Chen et al. [27] also showed the benefit of employing both hydrogen reduction and WE from Martian soil through the optimization of technology selection in terms of crewed Mars mission cost. Even though this study showed the economic benefit, they also noted the potential increase in the complexity of its development, deployment, and operation.
@@ 2.2 Mass and power estimation for ISRU
As an example of relatively low-fidelity modeling, Chen et al. [27] developed a database of subsystem-level specific masses and powers based on literature, such as Santiago-Maldonado and Linne [33], and integrated it into an Earth-Moon-Mars space resources logistics network. This approach analyzed the effect of ISRU deployment location (i.e., the Moon or a near-Earth object) regarding system mass and monetary mission cost.
While these models can capture the trend in mass and power for different architectures, they often overlook the detailed ConOps, potentially missing some key subsystems and components to be considered. Recently, more detailed models considering ConOps carefully have been developed. Linne et al. [7] modeled in detail the extraction of oxygen via CR of dry regolith to identify the feasibility of deploying such a plant considering the capacity of a lunar lander. Guerrero-Gonzalez and Zabel [32] also accounted for metal processing by comparing three different ISRU techniques, hydrogen reduction, molten regolith electrolysis, and molten salt electrolysis, and estimating the mass and required power of each subsystem. Kiewiet et al. [22] modeled WE from icy regolith via three different heating methods and conducted a trade-off study. Kleinhenz and Paz [23] compared the system mass and power of two different architectures: the extraction of both H2 and O2 from water ice in the lunar southern polar region and solely O2 based on the work developed by Linne et al. [7].
@@ 2.3 Recognizing uncertainty in ISRU
None of these studies have compared different ISRU technologies considering the effect of the uncertain lunar environment and ISRU operations. Due to the large difference between dry and icy regolith processing, potential benefits and risks can be easily overlooked when ignoring the uncertainty. Inspired by Cilliers et al. [30], this paper models the distributions of uncertain parameters and examines their effects on plant performance indicators such as regolith excavation rate.
@@ 3 Hybrid lunar ISRU plant concept
This section first proposes two different hybrid lunar ISRU architectures combining two extraction technologies. Technologies selected for further analyses and the detailed ConOps are explained in Sec. 3.2 and 3.3, respectively.
@@ 3.1 Overview and assumptions
Figure 1 depicts rather conventional single-technology ISRU architectures as well as hybrid architectures proposed in this paper. A typical dry regolith processing ISRU architecture, such as CR or hydrogen reduction, has a regolith excavation site, and a disposal site as well as a processing site (Fig. 1(a)). For WE from icy regolith, among various different kinds of architectures, an architecture suggested by Kleinhenz and Paz [23] is depicted in Fig. 1(b). In this architecture, water tankers transport water from a PSR to a PEL to further conduct electrolysis and liquefaction. For other WE technologies, please refer to Section 3.2.
Two different types of hybrid production plants are considered in this paper. The first architecture considered is a Parallel Hybrid (PH) plant that excavates and extracts resources from both PEL and PSR sites (Fig. 1(c)). In the PEL, dry regolith is excavated and processed to extract oxygen. Simultaneously in the PSR, water is extracted from icy regolith. Following the WE plant design proposed by Kleinhenz and Paz [23], the extracted water is transported from the PSR to the PEL by a mobile water transportation solution, where the water is then electrolyzed into O2 and H2. Thermochemical extraction processes, such as carbothermal or hydrogen reduction of dry regolith, produce water as an intermediate product, and therefore, water electrolysis and storage of oxygen and hydrogen can be shared between dry and icy regolith processing.
The second architecture considered is a Series Hybrid (SH), where dry regolith tailings from the WE of icy regolith are further processed directly in the PSR (Fig. 1(d)). Similar to the WE and PH architectures, the extracted and produced water is transported to a PEL for further processing. The detailed ConOps of both hybrid plants are further described in Section 3.3.
(a) Typical dry regolith processing architecture.
(b) Direct water extraction architecture based on Kleinhenz and Paz [23].
(c) Parallel hybrid architecture. Dry regolith and icy regolith are processed simultaneously in different sites.
(d) Series hybrid architecture. Dry regolith from WE process is further processed in a PSR.
Figure 1: ISRU production plant architectures considered in this paper.
Alternatively, transporting regolith excavated in a PSR, as well as water from WE, to a PEL for CR and further process is possible. This approach can reduce the number of excavators compared to the PH architecture. However, carrying excavated regolith from the PSR to the PEL requires a regolith transporting solution, such as a tanker. The results from Kleinhenz and Paz [23] indicate the mass of such tankers can be much more than the saved mass from excavators, and therefore this option is not considered further.
@@
Although identifying a specific site within the lunar South Pole was out of the scope of this work, previous studies [23, 42, 43, 44] have identified traverse distances of around 0.65-8.5 km between potential human landing system sites, permanently shadowed craters where acceptable water ice deposits might be found, and ridges with prolonged sunlight availability. A preliminary analysis demonstrated that mass and power budgets were not sensitive to changes in the traverse distance. Therefore, an average traverse distance of 5 km between PSRs and PELs is employed in this study as depicted in Fig. 1.
As a baseline case scenario, the ISRU plant is assumed to produce 10 t of O2 and 1.25 t of H2 per year. This 8:1 ratio corresponds to the stoichiometric mass ratio of water to maximize the use of local resources for various purposes. It is worth noting that, for hydrolox lunar ascent vehicle propellant, the fuel mixture mass ratio is generally close to 6:1. Some past studies employed this 6:1 ratio or a similar ratio as their required products ratio allowing an excess of oxygen [23, 45].
The integrated hybrid production plant model in this paper includes all the necessary ISRU subsystems, from regolith excavation to cryo-storage of the produced O2 and H2 (see Section 4 for a detailed explanation of the models), except for power and water purification subsystems. Power system is not considered in this paper for the sake of simplicity of analysis. Choosing an appropriate power generation system (e.g., photovoltaic, nuclear fission or radioisotope thermoelectric generator) as well as an energy storage system (e.g., battery or fuel cell) requires a large number of assumptions and can add significant complexity in the study. Furthermore, these systems can be shared with crewed a lunar base as well, which is out of the scope of this paper. The interested reader is referred to Refs. [27, 46] for power and energy storage systems on the Moon. It should be noted that power systems can play a critical role in the mass estimation of ISRU systems, and it should be assessed in the future. The infrastructure and operations to further utilize these products are not addressed in this work.
@@
Table 1 collects the global assumptions considered in this study, including the radiative outer space and average surface temperatures for each site, which are used for the heat loss and radiative heat exchange calculations, and the dry and icy regolith density dependencies on depth and water ice content. Individual assumptions are described for each model in Section 4. Moreover, Section 6 discusses the values of the following uncertain parameters considered in this study: feedstock particle size, silicate and water ice content in the regolith, WE efficiency, and ISRU plant operational availability.
@@
Avg. silica content 45 wt%
Avg. water ice content 4.2 wt%
Dry regolith density1 $\rho_{\mathrm{dry}}(z~{}[\mathrm{m}])=1800-700\cdot\exp(-z/0.06)~{}[\mathrm{kg}/\mathrm{m}^{3}]$
Icy regolith density2 $\rho_{\mathrm{icy}}(w_{\mathrm{ice}}~{}[\mathrm{wt\%}])=\rho_{\mathrm{dry}}\cdot(1+w_{\mathrm{ice}})~{}[\mathrm{kg}/\mathrm{m}^{3}]$
1 The regolith depth in meters is denoted by $z$, with $z=0~{}[m]$ at the lunar surface. Based on Hayne et al. [48].
2 The icy regolith density is calculated assuming that water ice mass fraction ($w_{\mathrm{ice}}$) accumulates uniformly within the regolith pores.
@@ 3.2 Technology selection
Over the last decades, numerous strategies for extracting oxygen from dry lunar regolith have been proposed in the literature [25, 4, 49]. From these strategies, ilmenite reduction by hydrogen followed by the electrolysis of water and carbon-based reductants (e.g., CO, CH4, C) has been considered as feasible and viable options [49]. CR of partially molten regolith, with water electrolysis and CH4-reforming, the direct electrochemical reduction of molten regolith, the electrolysis in a bath of molten salts, or vacuum thermal decomposition of lunar soils, have also been extensively researched [4].
CR of partially molten regolith is chosen as O2 extraction technology in this study due to its high yield, as it can reduce most regolith minerals, including the most prevalent silicates. Hydrogen reduction is excluded from further consideration since hydrogen is mostly limited to reducing iron-bearing minerals such as ilmenite, which are expected to only present in about 0.5 wt.of the lunar Sothern polar highland regolith [50]. As demonstrated by Guerrero-Gonzalez and Zabel [32], the low ilmenite content significantly increases the total hardware mass of the plant, highlighting its ineffectiveness as an oxygen extraction technology in these polar regions.
@@ 3.3.1 Parallel hybrid architecture
Six systems are involved in the overall concept of operations of the PH ISRU production architecture: (1) a dry-regolith mining (excavation and disposal) system, (2) a dry-regolith processing plant, (3) an H2/O2 production and storage plant, (4) an icy-regolith mining (excavation and disposal) system, (5) an icy-regolith processing plant, and (6) a water transportation system between both ISRU plants, as depicted in Fig. 1(c).
At a PEL, an excavator delivers dry regolith to the processing plant to extract oxygen. To avoid disturbances, the excavation site is located 100 m away from the ISRU facility [7]. After delivering a fresh regolith batch, the same vehicle is loaded with ISRU-produced tailings and drives first to a disposal site, also located 100 m away, before going back to the excavation site. The dry-regolith excavator recharges at the dry-regolith processing plant.
At a PSR, a similar scenario is defined. Another excavator delivers icy regolith to the icy-regolith processing plant, collects dry processed regolith, and dumps it at a disposal site before going back to the icy-regolith excavation site. The three locations are also positioned 100 m away from each other. The number of excavators for the dry- and icy-regolith is calculated based on the processing times, travel times, and regolith throughputs needed. The icy-regolith excavator only recharges at the icy-regolith processing plant without leaving the PSR.